OpenStudy (anonymous):

Tammy is at the dentist's office waiting on her appointment. She notices that the 6-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? Part 4: What is the coordinate point associated with this radian measure?

OpenStudy (welshfella):

use the formula arc length = radius * angle in radians L = rC

OpenStudy (anonymous):

That's what I tried but then I think I am now just over complicating it. So it would be 6(5π)? 30π?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i think you have it backwards the radian measure of the angle is the arc length divided by the radius

OpenStudy (misty1212):

if the arc length is \(5\pi\) and the radius is \(6\) the the measure of the angle is \[\frac{5\pi}{6}\]

OpenStudy (anonymous):

Oh, I was trying to solve for arc length. Duh. I see it now. Thank you so much!! So with that I can easily find the coordinates based on the unit circle.

OpenStudy (misty1212):

i hope so yes